A founder owns 100% of their startup. They are offered an equity investment by a VC investor, accepts, and eventually, undergoes one more round of financing, with a new VC investor. The financing events are as follows: VC investor 1 steps in with $0.5 million at a pre-money value of $2 million; later, VC investor 2 contributes $3 million at a pre-money of $7 million; After the second round of investment, what is the worth in stock of the founder, of VC1, and of VC2? What percentage of the company does each own?
Answer #1

All financials are in $ mn.

VC1 investment:

  • Pre money = 2
  • Investment = 0.5
  • Post money = Pre money + investment = 2 + 0.5 = 2.5
  • VC1's ownership = Investment / Post money = 0.5 / 2.5 = 20%
  • Founder's share = 1 - 20% = 80%

VC2 investment:

  • Pre money = 7
  • Investment = 3
  • Post money = Pre money + investment = 7 + 3 = 10

Answers to the questions follow sequentially:

  • Worth in stock of founder = Founder's share prior to VC2 investment x Pre money = 80% x 7 = 5.6 = $ 5.6 million
  • Worth in stock of VC1 = VC1's share prior to VC2 investment x Pre money = 20% x 7 = 1.4 = $ 1.4 million
  • Worth in stock of VC2 = Investment amount = 3 = $ 3 million
  • Founder's ownership = Worth in stock of founder / Post money = 5.6/10 = 56%
  • VC1's ownership = Worth in stock of VC1 / Post money = 1.4 / 10 = 14%
  • VC2's ownership = Investment / Post money = 3 / 10 = 30%